40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Sep 8 13:20:07 2011, steve, Update, PEM, vac envelope at atm for 36 days vent70.jpg
    Reply  Fri Jun 8 15:24:08 2012, steve, Update, VAC, RGA scan at day 276 and pressure plots 40m71pdM276d.pngpd71m276d.pngpd71d276.png
Message ID: 6790     Entry time: Fri Jun 8 15:24:08 2012     In reply to: 5366
Author: steve 
Type: Update 
Category: VAC 
Subject: RGA scan at day 276 and pressure plots 

Quote:

We tried not to open chambers above 10,000 particles of 0.5 micron cf/min

New items going in:                       2 rasor beam traps, 5 badly oxidized old silver plated  setscrew with spring loaded tips......to be replaced in the future, viton tips for eq screws....some are lose, gold plated allen wrench installed at ITMX bottom, reglued magnet on ITMX

Bad hardware things found:          nylon ball "locking elements" on OSEM locking set screws with screwdriver  slot, lose  1064 nm filter on OSEM pd

 40m - pumpdown 71- maglev pumping speed - at day 276........all normal

Attachment 1: 40m71pdM276d.png  45 kB  | Hide | Hide all
40m71pdM276d.png
Attachment 2: pd71m276d.png  30 kB  | Hide | Hide all
pd71m276d.png
Attachment 3: pd71d276.png  57 kB  | Hide | Hide all
pd71d276.png
ELOG V3.1.3-